# Asymptotic behaviour of the Hitchin metric on the moduli space of Higgs bundles

### v2

@@ 2.2.3 Deformation complex in the case where $\Sigma_{\theta}$ is smooth
Let $\tau$ be a holomorphic section of $\mathop{\rm End}\nolimits(E)\otimes K_{X}$ such that $\tau$ is mapped to 00 in $\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits(\theta))\simeq\pi_{\ast}(K_{\Sigma_{\theta}}^{1})$. There exists a holomorphic section $s$ of $\mathop{\rm End}\nolimits(E)$ such that $\tau=[\theta,s]$. Let $\tau^{\lor}$ (resp. $s^{\lor}$) denote the adjoint of $\tau$ (resp. $s$) with respect to $C$. Because $\theta^{\lor}=\theta$, we have $\tau^{\lor}=-[\theta,s^{\lor}]$. If $\tau=\pm\tau^{\lor}$, we have $\tau=\frac{1}{2}[\theta,s\mp s^{\lor}]$. We also note that any holomorphic section of $\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)\otimes K_{X}$ is mapped to 00 in $\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits(\theta))$. Then, we obtain $\mathcal{H}^{1}\bigl{(}\mathop{\rm Def}\nolimits(E,\theta;C)\bigr{)}\simeq\pi_{\ast}K_{\Sigma_{\theta}}$.
@@ 2.2.5 Locally defined non-degenerate symmetric pairings
[…]
@@
The restriction of $C_{Q|P}$ to $\mathcal{F}^{j}(E_{Q|P})\otimes\mathcal{F}^{k}(E_{Q|P})$ is 00 if $k+j>r_{Q}-1$. The induced pairings $\mathop{\rm Gr}\nolimits_{\mathcal{F}}^{j}(E_{Q|P})\otimes\mathop{\rm Gr}\nolimits_{\mathcal{F}}^{r_{Q}-1-j}(E_{Q|P})\longrightarrow{\mathbb{C}}$ are perfect. In particular, $C_{Q|P}$ is non-degenerate.
@@
There uniquely exists $g_{P}\in\mathop{\rm End}\nolimits(E_{|P})$ such that (i) $[f_{P|P},g_{P}]=\mathop{\rm Res}\nolimits_{P}(F_{\nu})$, (ii) $g_{P}$ is anti-self-adjoint with respect to $C_{P|P}$.
@@ 2.2.6 Dolbeault resolutions
The Cauchy operator of $E$ is denoted by ¯$\overline{\partial}_{E}$. The induced Cauchy operators of $\mathop{\rm End}\nolimits(E)$ is also denoted by ¯$\overline{\partial}_{E}$.
@@
Suppose that $\tau\in A^{1}(\mathop{\rm End}\nolimits(E))$ is a $1$-cocycle, i.e., (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\tau=0$. Then, $\tau$ induces a holomorphic section of $\mathcal{H}^{1}(\mathop{\rm Def}\nolimits(E,\theta))\simeq\pi_{\ast}K_{\Sigma_{\theta}}$. The induced holomorphic section is denoted by $\rho(\tau)$.
@@
Let $\tau\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta;C))$ such that (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\tau=0$ and that $\rho(\tau)=0$. Then, there exists a unique section $\sigma$ of $A^{0}(\mathop{\rm Def}\nolimits(E,\theta;C))$ such that (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma=\tau$ and that $\mathop{\rm Supp}\nolimits(\sigma)\subset\mathop{\rm Supp}\nolimits(\tau)$.
@@
Let $\tau\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta;C))$ such that (i) (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\tau=0$, (ii) $\mathop{\rm Supp}\nolimits\tau\subset X^{\circ}$, (iii) $\rho(\tau)=0$. Then, there exists a unique section $\sigma$ of $A^{0}(\mathop{\rm Def}\nolimits(E,\theta;C))$ such that (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma=\tau$ and that $\mathop{\rm Supp}\nolimits(\sigma)\subset\mathop{\rm Supp}\nolimits(\tau)$.
Proof   By the conditions, there exists a unique $C^{\infty}$-section $\sigma$ of $\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)$ such that (i) $\mathop{\rm ad}\nolimits(\theta)(\sigma)=\tau^{1,0}$, (ii) $\mathop{\rm Supp}\nolimits(\sigma)\subset\mathop{\rm Supp}\nolimits(\tau^{1,0})$. Because $\mathop{\rm ad}\nolimits(\theta):\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)\to\mathop{\rm End}\nolimits^{\mathop{\rm sym}\nolimits}(E;C)\otimes K_{X}$ is a monomorphism on $X^{\circ}$, it is easy to check that $\tau^{0,1}-\overline{\partial}_{E}\sigma=0$.
Let $U$ be any non-compact connected open subset of $X$. Let $\tau\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta;C)_{|U})$ such that (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\tau=0$ and that $\rho(\tau)=0$.
@@
There exists a unique section $\sigma$ of $A^{0}(\mathop{\rm Def}\nolimits(E,\theta;C)_{|U})$ such that (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma=\tau$ and that $\mathop{\rm Supp}\nolimits\sigma\subset\mathop{\rm Supp}\nolimits\tau$.
Proof   Because $U$ is non-compact, there exists a $C^{\infty}$-section $\sigma_{1}$ of $\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)$ such that ¯$\overline{\partial}_{E}\sigma_{1}=\tau^{0,1}$. We set $\tau_{1}=\tau-(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma_{1}$. Then, $\tau_{1}$ is a holomorphic section of $\mathop{\rm End}\nolimits^{\mathop{\rm sym}\nolimits}(E;C)\otimes K_{X}$ which is mapped to 00 in $\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits(\theta))$. Because $U$ is non-compact, there exists a unique holomorphic section $\sigma_{2}$ of $\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)$ such that $\mathop{\rm ad}\nolimits(\theta)(\sigma_{2})=\tau_{1}$. We set $\sigma=\sigma_{1}+\sigma_{2}$, and then (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma=\tau$ holds. By using $\mathop{\rm ad}\nolimits(\theta)(\sigma)=\tau^{1,0}$ we obtain the uniqueness of $\sigma$ and $\mathop{\rm Supp}\nolimits\sigma\subset\mathop{\rm Supp}\nolimits\tau$.
Let $N$ be an open neighbourhood of $D$ such that each connected component is non-compact. By applying Lemma 2.14 to $\tau_{|N}$, there exists $\sigma_{N}\in A^{0}(\mathop{\rm Def}\nolimits(E,\theta;C)_{|N})$ such that (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma_{N}=\tau_{|N}$. Let $N^{\prime}$ be a neighbourhood of $D$ in $N$ such that the closure of $N^{\prime}$ in $X$ is contained in $N$. Let $\chi:X\to[0,1]$ be a $C^{\infty}$-function such that $\chi=1$ on $N^{\prime}$ and $\chi=0$ on $X\setminus N$. We set $\tau_{1}=\tau-(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)(\chi\sigma_{N})$. Because $\mathop{\rm Supp}\nolimits(\tau_{1})\subset X^{\circ}$, there exists $\sigma_{1}\in A^{0}(\mathop{\rm Def}\nolimits(E,\theta;C))$ such that (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma_{1}=\tau_{1}$ and that $\mathop{\rm Supp}\nolimits(\sigma_{1})\subset\mathop{\rm Supp}\nolimits(\tau_{1})$. We set $\sigma=\chi_{N}\sigma_{N}+\sigma_{1}$. Then, we have (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma=\tau$. By using $\mathop{\rm ad}\nolimits(\theta)(\sigma)=\tau^{1,0}$, we obtain the uniqueness of $\sigma$ and $\mathop{\rm Supp}\nolimits\sigma\subset\mathop{\rm Supp}\nolimits\tau$.
@@ 2.2.7 Appendix
  • If $\nu$ is a holomorphic function, then $F_{\nu}$ satisfies (¯$(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)F_{\nu}=0$. In particular, $(\partial_{E,h}+\mathop{\rm ad}\nolimits\theta_{h}^{\dagger})F_{\nu}$ is a harmonic $1$-form of a harmonic bundle $(E,\theta,h)$. (See Proposition 2.26 below.)
  • If $\nu$ is a holomorphic $(1,0)$-form, then ¯$\overline{\partial}_{E}F_{\nu}=0$. If $\nu$ is a $(0,1)$-form, then $\mathop{\rm ad}\nolimits\theta(F_{\nu})=0$.
@@ 2.3.1 Infinitesimal deformations of a Higgs bundle
Let $(\widetilde{E},\widetilde{\theta})$ be an infinitesimal deformation of $(E,\theta)$. There exist isomorphisms $a_{i}:q^{\ast}(E_{|U_{i}})\simeq\widetilde{E}_{|U_{i}}$. Let $\psi_{i,j}$ denote the composition of the following isomorphisms:
@@
There exists $\rho_{i,j}\in H^{0}(U_{i}\cap U_{j},\mathop{\rm End}\nolimits(E))$ such that $\psi_{i,j}=\mathop{\rm id}\nolimits+\epsilon\rho_{i,j}$. We have $\rho_{j,k}-\rho_{i,k}+\rho_{i,j}=0$ on $U_{i}\cap U_{j}\cap U_{k}$. We also obtain $\kappa_{i}\in H^{0}(U_{i},\mathop{\rm End}\nolimits(E)\otimes K_{U_{i}})$ such that $a_{i}^{-1}\circ\widetilde{\theta}_{|U_{i}^{[1]}}\circ a_{i}=q_{U_{i}}^{\ast}(\theta)+\epsilon\kappa_{i}$. We have $\kappa_{j}-\kappa_{i}-\mathop{\rm ad}\nolimits(\theta)\rho_{i,j}=0$ on $U_{i}\cap U_{j}$. Then, the tuple
($(\{\rho_{i,j}\},\{\kappa_{i}\})\in\bigoplus_{i,j}H^{0}(U_{i}\cap U_{j},\mathop{\rm End}\nolimits(E))\oplus\bigoplus_{i}H^{0}(U_{i},\mathop{\rm End}\nolimits(E)\otimes K_{U_{i}})$
@@
Let ($(\{\widetilde{\rho}_{i,j}\},\{\widetilde{\kappa}_{i}\})$ is a $1$-cocycle induced by different trivializations $\widetilde{a}_{i}$. There exists $b_{i}\in H^{0}(U_{i},\mathop{\rm End}\nolimits(E))$ such that $\widetilde{a}_{i}=a_{i}\circ(1+\epsilon b_{i})$. Then, we have $\widetilde{\rho}_{i,j}=\rho_{i,j}+b_{j}-b_{i}$ on $U_{i}\cap U_{j}$ and $\widetilde{\kappa}_{i}=\kappa_{i}+\mathop{\rm ad}\nolimits(\theta)b_{i}$ on $U_{i}$. Hence, ($(\{\widetilde{\rho}_{i,j}\},\{\widetilde{\kappa}_{i}\})$ is cohomologous to ($(\{\rho_{i,j}\},\{\kappa_{i}\})$.

### v3

@@ 2.2.3 Deformation complex in the case where $\Sigma_{\theta}$ is smooth
Let $\tau$ be a holomorphic section of $\mathop{\rm End}\nolimits(E)\otimes K_{X}$ such that $\tau$ is mapped to 0 in $\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits(\theta))\simeq\pi_{\ast}(K_{\Sigma_{\theta}}^{1})$. There exists a holomorphic section $s$ of $\mathop{\rm End}\nolimits(E)$ such that $\tau=[\theta,s]$. Let $\tau^{\lor}$ (resp. $s^{\lor}$) denote the adjoint of $\tau$ (resp. $s$) with respect to $C$. Because $\theta^{\lor}=\theta$, we have $\tau^{\lor}=-[\theta,s^{\lor}]$. If $\tau=\pm\tau^{\lor}$, we have $\tau=\frac{1}{2}[\theta,s\mp s^{\lor}]$. We also note that any holomorphic section of $\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)\otimes K_{X}$ is mapped to 0 in $\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits(\theta))$. Then, we obtain $\mathcal{H}^{1}\bigl(\mathop{\rm Def}\nolimits(E,\theta;C)\bigr)\simeq\pi_{\ast}K_{\Sigma_{\theta}}$.
@@ 2.2.5 Locally defined non-degenerate symmetric pairings
$\begin{CD}E_{Q}\otimes E_{Q}@>{\pi_{\ast}(C_{L,Q})}>{}>\pi_{\ast}\Bigl(\mathcal{O}_{\Sigma_{Q}}\bigl(-(r_{Q}-1)Q\bigr)\Bigr)@>{a}>{}>\mathcal{O}_{X}.\end{CD}$
@@
The restriction of $C_{Q|P}$ to $\mathcal{F}^{j}(E_{Q|P})\otimes\mathcal{F}^{k}(E_{Q|P})$ is 0 if $k+j>r_{Q}-1$. The induced pairings $\mathop{\rm Gr}\nolimits_{\mathcal{F}}^{j}(E_{Q|P})\otimes\mathop{\rm Gr}\nolimits_{\mathcal{F}}^{r_{Q}-1-j}(E_{Q|P})\longrightarrow{\mathbb{C}}$ are perfect. In particular, $C_{Q|P}$ is non-degenerate.
@@
There uniquely exists $g_{P}\in\mathop{\rm End}\nolimits(E_{|P})$ such that (i) $f_{P|P},g_{P}]=\mathop{\rm Res}\nolimits_{P}(F_{\nu})$, (ii) $g_{P}$ is anti-self-adjoint with respect to $C_{P|P}$.
@@ 2.2.6 Dolbeault resolutions
The Cauchy operator of $E$ is denoted by $\overline{\partial}_{E}$. The induced Cauchy operators of $\mathop{\rm End}\nolimits(E)$ is also denoted by $\overline{\partial}_{E}$.
@@
Suppose that $\tau\in A^{1}(\mathop{\rm End}\nolimits(E))$ is a $1$-cocycle, i.e., $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\tau=0$. Then, $\tau$ induces a holomorphic section of $\mathcal{H}^{1}(\mathop{\rm Def}\nolimits(E,\theta))\simeq\pi_{\ast}K_{\Sigma_{\theta}}$. The induced holomorphic section is denoted by $\rho(\tau)$.
@@
Let $\tau\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta;C))$ such that $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\tau=0$ and that $\rho(\tau)=0$. Then, there exists a unique section $\sigma$ of $A^{0}(\mathop{\rm Def}\nolimits(E,\theta;C))$ such that $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma=\tau$ and that $\mathop{\rm Supp}\nolimits(\sigma)=\mathop{\rm Supp}\nolimits(\tau)$.
@@
Let $\tau\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta;C))$ such that (i) $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\tau=0$, (ii) $\mathop{\rm Supp}\nolimits\tau\subset X^{\circ}$, (iii) $\rho(\tau)=0$. Then, there exists a unique section $\sigma$ of $A^{0}(\mathop{\rm Def}\nolimits(E,\theta;C))$ such that $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma=\tau$ and that $\mathop{\rm Supp}\nolimits(\sigma)=\mathop{\rm Supp}\nolimits(\tau)$.
Proof  By the conditions, there exists a unique $C^{\infty}$-section $\sigma$ of $\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)$ such that (i) $\mathop{\rm ad}\nolimits(\theta)(\sigma)=\tau^{1,0}$, (ii) $\mathop{\rm Supp}\nolimits(\sigma)=\mathop{\rm Supp}\nolimits(\tau^{1,0})$. Because $\mathop{\rm ad}\nolimits(\theta):\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)\to\mathop{\rm End}\nolimits^{\mathop{\rm sym}\nolimits}(E;C)\otimes K_{X}$ is a monomorphism on $X^{\circ}$, it is easy to check that $\tau^{0,1}-\overline{\partial}_{E}\sigma=0$.
Let $U$ be any non-compact connected open subset of $X$. Let $\tau\in A^{1}(\mathop{\rm Def}\nolimits(E,\theta;C)_{|U})$ such that $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\tau=0$ and that $\rho(\tau)=0$.
@@
There exists a unique section $\sigma$ of $A^{0}(\mathop{\rm Def}\nolimits(E,\theta;C)_{|U})$ such that $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma=\tau$ and that $\mathop{\rm Supp}\nolimits\sigma=\mathop{\rm Supp}\nolimits\tau$.
Proof  Because $U$ is non-compact, there exists a $C^{\infty}$-section $\sigma_{1}$ of $\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)$ such that $\overline{\partial}_{E}\sigma_{1}=\tau^{0,1}$. We set $\tau_{1}=\tau-(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma_{1}$. Then, $\tau_{1}$ is a holomorphic section of $\mathop{\rm End}\nolimits^{\mathop{\rm sym}\nolimits}(E;C)\otimes K_{X}$ which is mapped to 0 in $\mathop{\rm Cok}\nolimits(\mathop{\rm ad}\nolimits(\theta))$. Because $U$ is non-compact, there exists a unique holomorphic section $\sigma_{2}$ of $\mathop{\rm End}\nolimits^{\mathop{\rm asym}\nolimits}(E;C)$ such that $\mathop{\rm ad}\nolimits(\theta)(\sigma_{2})=\tau_{1}$. We set $\sigma=\sigma_{1}+\sigma_{2}$, and then $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma=\tau$ holds. By using $\mathop{\rm ad}\nolimits(\theta)(\sigma)=\tau^{1,0}$ we obtain the uniqueness of $\sigma$ and $\mathop{\rm Supp}\nolimits\sigma=\mathop{\rm Supp}\nolimits\tau$.
Let $N$ be an open neighbourhood of $D$ such that each connected component is non-compact. By applying Lemma 2.14 to $\tau_{|N}$, there exists $\sigma_{N}\in A^{0}(\mathop{\rm Def}\nolimits(E,\theta;C)_{|N})$ such that $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma_{N}=\tau_{|N}$. Let $N^{\prime}$ be a neighbourhood of $D$ in $N$ such that the closure of $N^{\prime}$ in $X$ is contained in $N$. Let $\chi:X\to[0,1]$ be a $C^{\infty}$-function such that $\chi=1$ on $N^{\prime}$ and $\chi=0$ on $X\setminus N$. We set $\tau_{1}=\tau-(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)(\chi\sigma_{N})$. Because $\mathop{\rm Supp}\nolimits(\tau_{1})\subset X^{\circ}$, there exists $\sigma_{1}\in A^{0}(\mathop{\rm Def}\nolimits(E,\theta;C))$ such that $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma_{1}=\tau_{1}$ and that $\mathop{\rm Supp}\nolimits(\sigma_{1})\subset\mathop{\rm Supp}\nolimits(\tau_{1})$. We set $\sigma=\chi_{N}\sigma_{N}+\sigma_{1}$. Then, we have $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)\sigma=\tau$. By using $\mathop{\rm ad}\nolimits(\theta)(\sigma)=\tau^{1,0}$, we obtain the uniqueness of $\sigma$ and $\mathop{\rm Supp}\nolimits\sigma=\mathop{\rm Supp}\nolimits\tau$.
@@ 2.2.7 Appendix
  • If $\nu$ is a holomorphic function, then $F_{\nu}$ satisfies $(\overline{\partial}_{E}+\mathop{\rm ad}\nolimits\theta)F_{\nu}=0$. In particular, $(\partial_{E,h}+\mathop{\rm ad}\nolimits\theta_{h}^{\dagger})F_{\nu}$ is a harmonic $1$-form of a harmonic bundle $(E,\theta,h)$. (See Proposition 2.26 below.)
  • If $\nu$ is a holomorphic $(1,0)$-form, then $\overline{\partial}_{E}F_{\nu}=0$. If $\nu$ is a $(0,1)$-form, then $\mathop{\rm ad}\nolimits\theta(F_{\nu})=0$.
@@ 2.3.1 Infinitesimal deformations of a Higgs bundle
[…]
@@
There exist $\rho_{i,j}\in H^{0}(U_{i}\cap U_{j},\mathop{\rm End}\nolimits(E))$ such that $\psi_{i,j}=\mathop{\rm id}\nolimits+\epsilon\rho_{i,j}$. We have $\rho_{j,k}-\rho_{i,k}+\rho_{i,j}=0$ on $U_{i}\cap U_{j}\cap U_{k}$. We also obtain $\kappa_{i}\in H^{0}(U_{i},\mathop{\rm End}\nolimits(E)\otimes K_{U_{i}})$ such that $a_{i}^{-1}\circ\widetilde{\theta}_{|U_{i}^{[1]}}\circ a_{i}=q_{U_{i}}^{\ast}(\theta)+\epsilon\kappa_{i}$. We have $\kappa_{j}-\kappa_{i}-\mathop{\rm ad}\nolimits(\theta)\rho_{i,j}=0$ on $U_{i}\cap U_{j}$. Then, the tuple
$(\{\rho_{i,j}\},\{\kappa_{i}\})\in\bigoplus_{i,j}H^{0}(U_{i}\cap U_{j},\mathop{\rm End}\nolimits(E))\oplus\bigoplus_{i}H^{0}(U_{i},\mathop{\rm End}\nolimits(E)\otimes K_{U_{i}})$
@@
Let $(\{\widetilde{\rho}_{i,j}\},\{\widetilde{\kappa}_{i}\})$ is a $1$-cocycle induced by different trivializations $\widetilde{a}_{i}$. There exist $b_{i}\in H^{0}(U_{i},\mathop{\rm End}\nolimits(E))$ such that $\widetilde{a}_{i}=a_{i}\circ(1+\epsilon b_{i})$. Then, we have $\widetilde{\rho}_{i,j}=\rho_{i,j}+b_{j}-b_{i}$ on $U_{i}\cap U_{j}$ and $\widetilde{\kappa}_{i}=\kappa_{i}+\mathop{\rm ad}\nolimits(\theta)b_{i}$ on $U_{i}$. Hence, $(\{\widetilde{\rho}_{i,j}\},\{\widetilde{\kappa}_{i}\})$ is cohomologous to $(\{\rho_{i,j}\},\{\kappa_{i}\})$.